OpenStudy (jaedyn):

Grant has 2 over 3 cup of ice cream in a bowl. He then removes two 1 over 6 cups of ice cream to share with his little brother. Which of the following best explains the amount of ice cream left in Grant's bowl? 2 over 6, because 2 over 3 minus 2 times 1 over 6 equals 2 over 3 minus 3 over 6 2 over 6, because 2 over 3 minus 2 times 1 over 6 equals 2 over 3 minus 2 over 6 1 over 6, because 2 over 3 minus 2 times 1 over 6 equals 2 over 3 minus 3 over 6 1 over 6, because 2 over 3 minus 2 times 1 over 6 equals 2 over 3 minus 2 over 6

OpenStudy (taylor0402):

he has 2/3 he gives (1/6 * 2) = 1/3 2/3 - 1/3 = 1/3 he has 1/3 of a cup left

OpenStudy (taylor0402):

fraction 2 over 6, because fraction 2 over 3 minus 2 times fraction 1 over 6 equals fraction 2 over 3 minus fraction 2 over 6